Circular No. 6332

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE P. Garnavich and P. Challis, Harvard-Smithsonian Center for Astro-

physics, on behalf of the High-Z Supernova Search Team, report the discovery of seven supernovae on CCD images taken at the Victor Blanco 4-m telescope at Cerro Tololo. Each supernova is within ∼ 1′′ of its host galaxy’s nucleus. Spectra were obtained by B. Leibundgut on Feb. 22 and 23 using the 3.6-m European Southern Observatory telescope. Following are given for each supernova the dates of their discovery frames, position, R magnitude, redshift, supernova type, and observing group (1 = N. Suntzeff, A. Clocchiatti, and B. Schmidt; 2 = 1 + E. Turner and A. Athey):

SN  1996 UT  α_{2000}  δ_{2000}  R  z type G
1996E Feb. 14 10^{h}10^{m}20.^{s}28 −12^{o}33′09′′.2 22.2 0.43 Ia (1)
1996F Feb. 14 10 50 03.62 − 9 22 30.2 22.5 0.44 II? (1)
1996G Feb. 14 12 28 53.84 − 0 04 44.6 23.3 0.41 II? (1)
1996H Feb. 14 12 28 52.78 − 0 05 41.0 22.8 0.62 Ia (1)
1996I Feb. 14 12 00 39.43 − 0 16 02.4 23.0 0.57 Ia (1)
1996J Feb. 20 10 10 33.21 −12 37 05.2 22.5 0.30 Ia (2)
1996K Feb. 20 8 24 43.49 − 0 21 06.2 23  0.38 Ia (2)

Narrow emission lines from the host galaxies appear in the spectra of supernovae 1996E, 1996F, 1996G, and 1996H. Redshifts for the other supernovae are based on correlating broad features with spectra of nearby type-Ia supernovae. Supernovae 1996F and 1996G each show a featureless blue continuum. Finder charts may be requested by sending electronic mail

to pgarnavich@cfa.harvard.edu.

COMET C/1996 B2 (HYAKUTAKE) R. M. West, European Southern Observatory (ESO), writes: “Two

spectra (400–800 nm; resolution 1.2 nm; exposures 10 min each) were obtained by K. Simon and C. Lidman (ESO) with the Danish 1.54-m telescope (+ DFOSC) on Feb. 29.3 UT and reduced by S. Benetti (ESO). The spectrum is typical for a comet at r = 1.5 AU. The strongest features are the Swan C_{2} (0,0) 516-nm, (1,0) 474-nm, and (0,1) 564-nm bands. Others identified are the C_{3} complex at 405 nm; CN (0,1) at 422 nm; [O I] at 630 and 636 nm; NH_{2} (10,0), (9,0), (8,0), and (7,0) at 570–662 nm. With the possible exception of CO^{+} at 427 nm, no ionic lines are present in these spectra.”

1996 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green